OpenStudy (anonymous):

sqareroot of 200 a^6 b^7

OpenStudy (xishem):

\[\sqrt{200a^6b^7}=10a^3b^3 \sqrt{2b}\]

OpenStudy (anonymous):

would you put absolute signs to the odd exponents?

OpenStudy (anonymous):

can you show me how to do this?